A certain sports drink has a sugar concentration of 0.30M according to the label. If the container has 200ml of the drink, how many moles of sugar would you consume when drinking 50ml of the drink?
Answer #1

Moles of sugar consumed = 0.015 moles.

EXPLANATION: Molarity is the number of moles of solute present in one liter of solution. The number of moles of solute present in a particular volume of solution with a given molarity is determine by the following relation:

Moles of solute present = molarity \times volume of solution in L

We have, molarity = 0.30 M = 0.30 mol/L and volume = 50 mL = (50/1000) L = 0.050 L

Now, moles of sugar consumed when drinking 50 mL of the drink = (0.30 mol/L) \times (0.050 L) = 0.015 moles.